# An Introduction to Single-Antenna Radio Astronomical Polarimetry

### v2

@@ G.2 Dual-Sideband Down-Conversion
Note that changing the sign of the 90-degree phase shift in equation (G3) changes the sign of $q(t)$, which results in complex conjugation of $z_{b}(t)$. This is equivalent to modulating $x(t)$ with a local oscillator that has frequency $\nu_{0}$, which shifts the negative half of the spectrum of $X(\nu)$ to baseband. As in the case of lower-sideband down-conversion, the resulting spectrum is complex conjugated and the frequency axis is negated.
@@ Appendix H Reflection as Turning Over the Receiver
Reflection can also be modeled as turning over the receiver, which is equivalent to rotating the reference frame depicted in Figure 1 by $180^{\circ}$ about a transverse axis in the $x$-$y$ plane. This negates the $z$ axis and results in a two-dimensional reflection in the $x$-$y$ plane through the rotation axis. Let $\hat{\underline{\mbox{$a$}}}=(\cos\alpha,\sin\alpha,0)$ define the three-dimensional rotation axis in the physical space of Figure 1 and use Rodrigues’ Rotation Formula to express the $3\times 3$ matrix for a 180 rotation about $\hat{\underline{\mbox{$a$}}}$.
@@
The polar decomposition of a two-dimensional reflection through the axis defined by $(\cos\alpha,\sin\alpha)$ yields a unitary matrix parameterized by $\hat{\underline{\mbox{$n$}}}=(\cos 2\alpha\,\sin 2\alpha,0)$,
@@
Congruence transformation by ${\bf W}(\alpha)$ rotates $\underline{\mbox{$S$}}$ by 180-deg about the $\hat{\underline{\mbox{$n$}}}$ axis defined by $\alpha$, which negates both the ellipticity angle $\chi$ and the position angle $\psi$.
In a basis defined by linearly-polarized receptors, the nominal axis of symmetry has a position angle of $\pi/4$ radians. The reflection defined by ${\bf W}(\pi/4)=\mbox{$\sigma$}_{2}$ swaps $e_{x}$ and $e_{y}$. With reference to equations (18) through (20), swapping $e_{x}$ and $e_{y}$ negates Stokes Q and V. This negation is equivalent to a $\pm 180^{\circ}$ rotation of the Stokes polarization vector about $\hat{\underline{\mbox{$n$}}}=(0,1,0)$, which is also the result of congruence transformation by ${\bf W}(\pi/4)$.
In a basis defined by circularly-polarized receptors, the nominal axis of symmetry has a position angle of 0 (zero), and the reflection ${\bf W}(0)=\mbox{$\sigma$}_{1}$ negates $e_{y}$. With reference to equations (18) through (20), negating either component of the electric field vector reverses the signs of both Stokes U and V, which is equivalent to the $\pm 180^{\circ}$ rotation of the Stokes polarization vector about $\hat{\underline{\mbox{$n$}}}=(1,0,0)$ under congruence transformation by ${\bf W}(0)$.
@@ Appendix I Fundamental Numerical Instability
In this measurement equation, $\underline{\underline{{\bf X}}}$ is the known constraining transformation, or matrix argument, and $\alpha$ is the set of model parameters that describe both $\underline{\underline{{\bf M}}}$ and $\underline{\underline{\mbox{$S$}}}$. The matrix argument $\underline{\underline{{\bf X}}}$ can be the projection between the receptors and the celestial sphere (Section 5.5), which can vary with time owing to the rotation of the Earth or mechanical rotation of the feed horn. Faraday rotation (Section 5.6) may also serve as the matrix argument under the assumption that the intrinsic Stokes parameters do not vary with radio frequency $\nu$, or are a known function of $\nu$ (e.g., Edwards & Stappers, 2004).
Both forms of numerical instability arise when solving equation (I1) through variation of $\alpha$. Model degeneracy occurs when unknown components of $\underline{\underline{{\bf M}}}$ commute with $\underline{\underline{{\bf X}}}$, and parameter collinearity arises when components of $\underline{\underline{\mbox{$S$}}}$ are eigenvectors of $\underline{\underline{{\bf X}}}$.
@@ I.1 Commuting with the Matrix Argument
Note that depolarization of Stokes V includes negation (c=1c=-1) and, as described in Section 5.1, no linear transformation of the electric field can negate the sign of Stokes V. Therefore, if the instrumental response is modeled using Jones matrices, then negation of Stokes V in the unknown $\underline{\underline{\mbox{$S$}}}$ can be compensated only by rotating the reference frame. However, a $\pm 180^{\circ}$ rotation that negates Stokes V must also negate the position angle and derived quantities such as the Faraday rotation measure. Therefore, the Stokes V sign ambiguity can be eliminated by observing a source for which the sign of the position angle and/or rotation measure is known.
@@
When modeling variations of the observed Stokes parameters as a function of a matrix argument with a variable axis of symmetry, it is no longer possible for an unknown component of $\underline{\underline{{\bf M}}}$ to commute with all values of $\underline{\underline{{\bf X}}}$, and the degeneracy is eliminated. However, as described in the following section, even when the axis of symmetry of $\underline{\underline{{\bf X}}}$ is variable, the model parameters can be highly collinear and cause numerical instability.
@@ I.2 Eigenvectors of the Matrix Argument
However, the Stokes vector, $\underline{\underline{\mbox{$V$}}}=[0,0,0,V]$ has a polarization vector that is parallel to the symmetry axis of $\underline{\underline{{\bf R}}}$ and perpendicular to the symmetry axis of $\underline{\underline{{\bf B}}}_{L}$. Therefore, it is an eigenvector of $\underline{\underline{{\bf P}}}$ with associated eigenvalue $\lambda=1$; i.e.,
@@
where $\underline{\underline{\mbox{$S$}}}_{n,L}=[I_{n},Q_{n},U_{n},0]$ and $\underline{\underline{\mbox{$S$}}}_{n,V}=[0,0,0,V_{n}]$, such that

### v4

@@ G.2 Dual-sideband Down-conversion
Note that changing the sign of the $90^{\circ}$ phase shift in Equation (G3) changes the sign of $q(t)$, which results in complex conjugation of $z_{b}(t)$. This is equivalent to modulating $x(t)$ with a local oscillator that has frequency $\nu_{0}$, which shifts the negative half of the spectrum of $X(\nu)$ to baseband. As in the case of lower-sideband down-conversion, the resulting spectrum is complex conjugated and the frequency axis is negated.
@@ Appendix H Reflection as Turning Over the Receiver
Reflection can also be modeled as turning over the receiver, which is equivalent to rotating the reference frame depicted in Figure 1 by $180^{\circ}$ about a transverse axis in the $x$-$y$ plane. This negates the $z$ axis and results in a two-dimensional reflection in the $x$-$y$ plane through the rotation axis. Let $\hat{\underline{\mbox{$a$}}}=(\cos\alpha,\sin\alpha,0)^{T}$ define the three-dimensional rotation axis in the physical space of Figure 1 and use Rodrigues’ Rotation Formula to express the $3\times 3$ matrix for a 180 rotation about $\hat{\underline{\mbox{$a$}}}$.
@@
The polar decomposition of a two-dimensional reflection through the axis defined by $(\cos\alpha,\sin\alpha)$ yields a unitary matrix parameterized by $\hat{\underline{\mbox{$n$}}}=(\cos 2\alpha\,\sin 2\alpha,0)^{T}$,
@@
Congruence transformation by ${\bf W}(\alpha)$ rotates $\underline{\mbox{$S$}}$ by $180^{\circ}$ about the $\hat{\underline{\mbox{$n$}}}$ axis defined by $\alpha$, which negates both the ellipticity angle $\chi$ and the position angle $\psi$.
In a basis defined by linearly-polarized receptors, the nominal axis of symmetry has a position angle of $45^{\circ}$. The reflection defined by ${\bf W}(\pi/4)=\mbox{$\sigma$}_{2}$ swaps $e_{x}$ and $e_{y}$. With reference to Equations (18) through (20), swapping $e_{x}$ and $e_{y}$ negates Stokes Q and V. This negation is equivalent to a $\pm 180^{\circ}$ rotation of the Stokes polarization vector about $\hat{\underline{\mbox{$n$}}}=(0,1,0)^{T}$, which is also the result of congruence transformation by ${\bf W}(\pi/4)$.
In the circular basis, the nominal axis of symmetry has a position angle of $0^{\circ}$, and ${\bf W}(0)=\mbox{$\sigma$}_{1}$ negates $e_{y}$. Negating either component of the electric field vector reverses the signs of both Stokes U and V, which is equivalent to the $\pm 180^{\circ}$ rotation of the Stokes polarization vector about $\hat{\underline{\mbox{$n$}}}=(1,0,0)^{T}$ under congruence transformation by ${\bf W}(0)$.
@@ Appendix I Fundamental Numerical Instability
In this measurement equation, $\underline{\underline{{\bf X}}}$ is the known constraining transformation, or matrix argument, and $\alpha$ is the set of model parameters that describe both $\underline{\underline{{\bf M}}}$ and $\underline{\underline{\mbox{$S$}}}$. The matrix argument $\underline{\underline{{\bf X}}}$ can be the projection between the receptors and the celestial sphere (Section V.5), which can vary with time owing to the rotation of the Earth or mechanical rotation of the feed horn. Faraday rotation (Section V.6) may also serve as the matrix argument under the assumption that the intrinsic Stokes parameters do not vary with radio frequency $\nu$, or are a known function of $\nu$ (e.g., Edwards & Stappers, 2004).
Both forms of numerical instability arise when solving Equation (I1) through variation of $\alpha$. Model degeneracy occurs when unknown components of $\underline{\underline{{\bf M}}}$ commute with $\underline{\underline{{\bf X}}}$, and parameter collinearity arises when unknown components of $\underline{\underline{\mbox{$S$}}}$ are eigenvectors of $\underline{\underline{{\bf X}}}$.
@@ I.1 Commuting with the Matrix Argument
Note that depolarization of Stokes V includes negation (c=1c=-1) and, as described in Section V.1, no linear transformation of the electric field can negate the sign of only Stokes V. For example, if the instrumental response is modeled using Jones matrices, then negation of Stokes V in the unknown $\underline{\underline{\mbox{$S$}}}$ can be compensated only by rotating the reference frame. However, a $\pm 180^{\circ}$ rotation that negates Stokes V must also negate the position angle and derived quantities such as the Faraday rotation measure. Therefore, the Stokes V sign ambiguity can be eliminated by observing a source for which the sign of the position angle and/or rotation measure is known.
@@
When modeling variations of the observed Stokes parameters as a function of a matrix argument with a variable axis of symmetry, it is no longer possible for an unknown component of $\underline{\underline{{\bf M}}}$ to commute with all values of $\underline{\underline{{\bf X}}}$, and the degeneracy is eliminated. However, as described in the following section, even when the axis of symmetry of $\underline{\underline{{\bf X}}}$ is variable, unknown model parameters can be highly collinear and cause numerical instability.
@@ I.2 Eigenvectors of the Matrix Argument
However, the Stokes vector, $\underline{\underline{\mbox{$V$}}}=[0,0,0,V]^{T}$ has a polarization vector that is parallel to the symmetry axis of $\underline{\underline{{\bf R}}}$ and perpendicular to the symmetry axis of $\underline{\underline{{\bf B}}}_{L}$. Therefore, it is an eigenvector of $\underline{\underline{{\bf P}}}$ with associated eigenvalue $\lambda=1$; i.e.,
@@
where $\underline{\underline{\mbox{$S$}}}_{n,L}=[I_{n},Q_{n},U_{n},0]^{T}$ and $\underline{\underline{\mbox{$S$}}}_{n,V}=[0,0,0,V_{n}]^{T}$, such that